OpenStudy (anonymous):

Solve 4 log12 2 + log12 x = log12 96 Choose one answer. a. x = 88 b. x = 80 c. x = 12 d. x = 6

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

your equation becomes \[4\times \log_{12}(2) + \log_{12} (x) - \log_{12}(96) =0\]

OpenStudy (anonymous):

\[\log_{12}(2^{4})+\log_{12}(x)+\log_{12}(96^{-1})=0 \]

OpenStudy (anonymous):

\[\log_{12}(16\times x \times 96^{-1})\]

OpenStudy (anonymous):

so the equation becomes simpler and you have to solve \[16\times x \times 96^{-1}=1\]

OpenStudy (anonymous):

the solution is \[x= 96/16=6\] the answer is D